# Gil Ayache*, Menachem Pirchi*, Aviv Navon, Aviv Shamsian, Gill Hetz, Joseph Keshet aiOla Research, Israel

### v1

@@ Abstract
Integrating named entity recognition (NER) with automatic speech recognition (ASR) can significantly enhance transcription accuracy and informativeness. In this paper, we introduce WhisperNER, a novel model that allows joint speech transcription and entity recognition. WhisperNER supports open-type NER, enabling recognition of diverse and evolving entities at inference. Building on recent advancements in open NER research, we augment a large synthetic dataset with synthetic speech samples. This allows us to train WhisperNER on a large number of examples with diverse NER tags. During training, the model is prompted with NER labels and optimized to output the transcribed utterance along with the corresponding tagged entities. To evaluate WhisperNER, we generate synthetic speech for commonly used NER benchmarks and annotate existing ASR datasets with open NER tags. Our experiments demonstrate that WhisperNER outperforms natural baselines on both out-of-domain open type NER and supervised finetuning.
@@ I Introduction
Recent advancements in both speech processing and natural language processing (NLP) have dramatically expanded the capabilities of automatic speech recognition (ASR) systems. These improvements, driven by large-scale models like Whisper [1], have significantly enhanced transcription accuracy. Meanwhile, the emergence of powerful NLP models, particularly large language models (LLMs), has led to major breakthroughs in tasks like language understanding, text generation, and named entity recognition (NER).
@@
Traditionally, speech-driven applications have relied on pipeline architectures, where ASR systems transcribe speech into text, which is then processed by NLP models for tasks such as sentiment analysis, question answering, and NER. While this approach has proven effective across a range of applications, it suffers from error accumulation — where transcription errors in the ASR stage propagate through the pipeline, reducing the performance of downstream NLP tasks. This challenge is particularly evident in complex tasks requiring high accuracy, such as NER.
@@
A growing area of interest is the integration of speech models into a variety of NLP tasks beyond transcription, allowing models to process spoken language directly and eliminate intermediate stages [2, 3, 4, 5]. For instance, spoken language understanding models have been developed to perform tasks like intent detection and NER [6, 2, 5] directly from speech, bypassing the need for separate ASR components. These end-to-end models have shown promising results by jointly optimizing for both speech and language understanding.
Figure 1: WhisperNER: A set of entity types is provided as prompt to the decoder, which is optimized to output the transcription along with the corresponding entity tags.
@@
TABLE I: Zero-shot, out-of-domain evaluation on three open type NER speech benchmarks. The Params column indicate the extra parameters of baseline methods w.r.t WhisperNER. We bold the best results and underline the second-best results.
@@
VoxPopuli-NER LibriSpeech-NER Fleurs-NER Average F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ Params $\downarrow$ GNER-T5-base $45.31$ $\mathbf{8.32}$ $35.68$ $\mathbf{5.51}$ $47.15$ $\underline{8.52}$ $42.72$ $\mathbf{7.45}$ $+248$M NuNER $52.18$ $\mathbf{8.32}$ $43.65$ $\mathbf{5.51}$ $53.11$ $\underline{8.52}$ $49.64$ $\mathbf{7.45}$ $+459$M GLiNER $53.55$ $\mathbf{8.32}$ $48.97$ $\mathbf{5.51}$ $\underline{54.35}$ $\underline{8.52}$ $52.29$ $\mathbf{7.45}$ $+459$M WhisperNER-BIO $\underline{55.79}$ $9.51$ $\underline{50.40}$ $5.95$ $\mathbf{54.42}$ $\mathbf{8.22}$ $\underline{53.53}$ $\underline{7.89}$- WhisperNER $\mathbf{57.24}$ $\underline{9.24}$ $\mathbf{50.84}$ $\underline{5.82}$ $53.50$ $10.26$ $\mathbf{53.86}$ $8.44$-
@@
Recently, open NER [7, 8, 9] has gained significant attention for its ability to generalize to new and unseen entities, offering a flexible alternative to traditional NER systems that are restricted by fixed entity types. [7] introduce a targeted distillation method, where student models are trained through instruction tuning. [8, 9] suggested using a tailored, small size transformer encoders, designed specifically for the NER problem, and showed that this models can outperform LLM based NER models. [10] propose a method for improving information extraction results through annotation guidelines, and [11] explored the importance of “negative instances”, i.e, non-entity spans, for improving contextual information and label boundaries. However, integrating open type NER directly into ASR systems remains an unexplored area.
In this paper, we propose WhisperNER, a unified model that simultaneously performs speech transcription and named entity recognition, based on the Whisper ASR model [1]. By integrating NER directly into the ASR process, WhisperNER eliminates the need for separate ASR and NER components, thereby reducing the risk of error propagation. The model supports open type NER, enabling it to recognize and tag a wide range of entities during inference, offering a more accurate and flexible approach to entity recognition from speech.
@@
To build WhisperNER, we leverage a large-scale synthetic dataset released from [9]. We augment a subset of the textual open NER dataset with corresponding synthetic audio samples. The final dataset contains $350$K samples and $\sim 1.8$M unique entity types. Our approach draws motivation from recent approaches which incorporates prior information and context into the decoding process by prompting the model on domain information [12] or keyword phrases [13, 14]. We train WhisperNER by prompting the model with NER type, for the task of joint transcription and NER tagging (See Figure 1). At inference time, WhisperNER can generalize to novel entity types not observed during training.
@@
This paper makes the following contributions: (i) Propose WhisperNER, a novel model for joint open NER and ASR; (ii) Introducing new speech and text NER and open-NER benchmarks (iii) Conduct extensive experiments to demonstrate the superiority of WhisperNER.
To encourage future research and reproducibility, we will make our source code, datasets, and models publicly available.
@@ II Our Method
In this section we describe our approach, WhisperNER, which extends the Whisper [1] ASR model to allow for joint open NER and speech recognition. The model is designed to output a tagged transcription sequence given an audio input, effectively performing both tasks simultaneously.
@@ II-A Model Architecture and Training
Let $\mathbf{x}$ denote the input audio features, which are fed into the encoder. The encoder produces a sequence of hidden states $\mathbf{h}=\text{Encoder}(\mathbf{x})$, which are then used to condition the generation process.
In our approach, we further condition the decoding process on a set of entity tags $\mathbf{t}=[t_{1},t_{2},\ldots,t_{k}]$ (where each $t_{i}$ represents a specific entity type such as person, location, or food). The decoder then generates the output sequence $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$, which includes both the transcribed text and the corresponding entity tags, see Figure 1.
Formally, the decoder generates each token $y_{t}$ based on the previous tokens and the encoder’s hidden states,
@@
To mark the entities in the output sequence, we explore two approaches. The second approach uses the uses the BIO tagging scheme, similar to the approach used in [11], i.e., ...walks(O) in(O) the(O) national(B-location) park(I-location). The second marks entities with start and end markers as follow <$t_{i}$>$y_{i_{1}}...y_{i_{n}}$<$t_{i}$>>, for example, <location>national park<location>> (see Figure 1). While the former may induce greater supervision and improve generalization [11], this approach increases the length of the generated sequence, hence it is potentially less compute and time efficient.
Our prompting based training procedure, allows WhisperNER to handle both traditional (i.e., predefined set of entities) and open-type NER tasks. By exposing the model to diverse entity types during training, we ensure that it can generalize well to new and unseen entities during inference, making it highly adaptable to a wide range of applications.
II-B Negative Sampling and Entity Type Dropout
In our experiments, we find that including negative samples, i.e., prompting the model with entity labels not presented in the speech utterance, to significantly enhance the model’s generalization performance, in line with findings in previous literature on open-type NER [8]. We explored various negative sampling approaches and chose to use the complete entity set from a negative example, randomly selected from the training set, which proved most effective.
Additionally, we applied random entity type dropout during training, as inspired by Sainz et al. [10]. This involves randomly eliminating a subset of entity types from the prompt and output sequence, compelling the model to focus on the remaining defined entity types. We found this approach to reduce entity hallucination risks during inference – where the model output entities not presented in the input prompt. Also, we incorporate random shuffling of entities in the input prompt to further promote generalization across diverse input sequences. For a detailed analysis of these methodologies and their impacts, see Section III-D.
@@ III Experiments
In this section, we evaluate WhisperNER against naturall, pipeline based baselines, using NER and open type NER speech datasets.
Datasets.  Our evaluation setup consists of open type NER benchmarks and commonly used NER benchmarks. For the open type NER benchmarks we utilize three frequently used speech datasets with diverse linguistic and contextual properties. These datasets include VoxPopuli [15], LibriSpeech [16], and Fleurs [17], each offering unique challenges and opportunities for evaluating NER models, particularly in open-type NER contexts. For the multilingual datasets, we focus specifically on the English subset of the data. As this ASR datasets do not include NER labels, we adopt an LLM based annotation approach, in line with the method described in [9]. We utilize the prompts derived approach to extract open type NER from the raw transcriptions of each dataset. See [9] for full details of the approach and the specific prompt. This approach results with a wide and diverse set of NER annotations, allowing us to test the capabilities of our approach to identify a broader range of entity types that are not confined to traditional categories; Thus, enhancing our understanding of the model’s generalization in real-world applications.
@@
This annotation scheme results with the following ASR-NER datasets: (i) VoxPopuli-NER: This dataset comprises $879$ samples. It is annotated with $2469$ unique entity types, providing a comprehensive basis for political and geographical entity recognition ; (ii) LibriSpeech-NER: Based on the LibriSpeech-clean dataset, with $1604$ samples derived from audio book readings, includes annotations for $3674$ unique entity types. This dataset is particularly valuable for testing the model’s ability to identify personal names and locations within narrative contexts; (iii) Fleurs-NER: Consists of $441$ samples, annotated for $1440$ unique entity types.
@@
In addition, we utilize commonly used, textual NER benchmarks [8, 9, 10, 11], namely MIT-Movie [18] and MIT-Restaurant [18]. These datasets consists of $\sim 14$K and $\sim 10$K, with $12$ and $8$ entity types, respectively. To enhance these text-based datasets for speech recognition, we generated corresponding audio data using the AWS Polly text-to-speech (TTS) service. To support future research on joint ASR and NER, we will release all datasets.
@@
Baselines.  We evaluate our WhisperNER againt natural two-stage baselines, in which we initially employ a Whisper large-v2 model for speech-to-text transcription, followed by a NER model. We constrain our evaluation to recent small-scale NER models (up to 0.5 billion parameters) to ensure competitiveness and maintain parity in terms of the number of parameters and computational requirements, aligning with our end-to-end approach. The baselines include: (i) GLiNER [8]: a transformer encoder based architecture with zero-shot NER capabilities. We use the large variant, GLiNER-L; (ii) NuNER [9]: since the original NuNER model is not suitable for the zero-shot NER task, we utilize a GLiNER based architecture model, trained with the synthetic dataset of [9], named NuNER zero-span.; (iii) GNER [11]: Different from the previous two benchmarks, this approach uses generative models specifically adopted for the NER task. We use the flan-t5-base variant; (iv) WhisperNER-BIO: Our approach with complete BIO annotations; (v) WhisperNER: Our approach with NER span boundary annotations (Figure 1).
@@
TABLE II: Zero-Shot Performance on the MIT NER Benchmarks.
Movie Restaurant F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ GNER-T5-base $39.98$ $9.92$ $\underline{30.45}$ $11.10$ NuNER $\mathbf{44.64}$ $9.92$ $\mathbf{34.04}$ $11.10$ GLiNER $\underline{40.44}$ $9.92$ $29.12$ $11.10$ WhisperNER-BIO $30.16$ 8.10 $26.67$ $\mathbf{8.44}$ WhisperNER $30.55$ $\underline{8.39}$ $28.22$ $\underline{9.64}$
Training details.  The WhisperNER models are trained using a subset of $350$K samples from the dataset proposed in [9]. For each example, we generate a corresponding synthetic audio sample. Unless stated otherwise, we train the WhisperNER models for $250$K steps, with learning rate $61e-$ 6 and linear decay learning rate scheduler. During training, we freeze the Whisper encoder and only modify the decoder.

### v2

# WhisperNER: Unified Open Named Entity and Speech Recognition

@@ Abstract
Integrating named entity recognition (NER) with automatic speech recognition (ASR) can significantly enhance transcription accuracy and enrich its content. We introduce WhisperNER, a novel model that facilitates joint speech transcription and entity recognition. WhisperNER supports open-type NER, enabling recognition of various entities during inference. Building on recent advancements in open NER research, we augment a large synthetic dataset with synthetic speech samples. This approach enables us to train WhisperNER on numerous examples with various NER tags. During training, the model is prompted with NER labels and optimized to produce the transcribed utterance alongside the corresponding tagged entities. For evaluation, we generate synthetic speech for commonly used NER benchmarks and annotate existing ASR datasets with open NER tags. Our experiments show that WhisperNER outperforms natural baselines in both out-of-domain open-type NER and supervised fine-tuning.
@@ I Introduction
Recent advancements in both speech processing and natural language processing (NLP) have dramatically expanded the capabilities of automatic speech recognition (ASR) systems. These improvements, driven by large-scale models like Whisper [1], have significantly enhanced transcription accuracy. Meanwhile, the emergence of powerful NLP models, particularly large language models (LLMs), has led to major breakthroughs in tasks like language understanding, text generation, and named entity recognition (NER). Traditionally, speech-driven applications have relied on pipeline architectures, where ASR systems transcribe speech into text, which is then processed by NLP models for tasks such as sentiment analysis, question answering, and NER. While this approach has proven effective across a range of applications, it suffers from error accumulation — where transcription errors in the ASR stage propagate through the pipeline, reducing the performance of downstream NLP tasks. This challenge is particularly evident in complex tasks requiring high accuracy, such as NER.
@@
A growing area of interest is the integration of speech models into a variety of NLP tasks beyond transcription, allowing models to process spoken language directly and eliminate intermediate stages [2, 3, 4, 5, 6]. For instance, spoken language understanding (SLU) models have been developed to perform tasks such as intent detection and NER [7, 8, 9] directly from speech, bypassing the need for separate ASR components. These end-to-end models have shown promising results by jointly optimizing speech and language understanding.
Figure 1: The architecture of WhisperNER. A set of entity types is provided as a prompt to the decoder. During training, we provide positive (green) and negative (red) entities. At inference time, the model can generalize to novel entity types not observed during training.
@@
TABLE I: Zero-shot, out-of-domain evaluation on three open-type NER speech benchmarks. The Params column indicates the extra parameters of baseline methods w.r.t WhisperNER (since baselines use a pipeline that consists of the Whisper model and the NLP model, while WhisperNER uses only the Whisper architecture). We bold the best results and underline the second-best results.
@@
VoxPopuli-NER LibriSpeech-NER Fleurs-NER Average F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ F1 $\uparrow$ WER $\downarrow$ Params $\downarrow$ GNER-T5-base $45.31$ $\mathbf{8.32}$ $35.68$ $\mathbf{5.51}$ $47.15$ $\underline{8.52}$ $42.72$ $\mathbf{7.45}$ +248+248M NuNER $52.18$ $\mathbf{8.32}$ $43.65$ $\mathbf{5.51}$ $53.11$ $\underline{8.52}$ $49.64$ $\mathbf{7.45}$ +459+459M GLiNER $53.55$ $\mathbf{8.32}$ $48.97$ $\mathbf{5.51}$ $\underline{54.35}$ $\underline{8.52}$ $52.29$ $\mathbf{7.45}$ +459+459M WhisperNER-BIO $\underline{55.79}$ $9.51$ $\underline{50.40}$ $5.95$ $\mathbf{54.42}$ $\mathbf{8.22}$ $\mathbf{53.53}$ $\underline{7.89}$ - WhisperNER $\mathbf{56.25}$ $\underline{9.22}$ $\mathbf{50.84}$ $\underline{5.82}$ $53.50$ $10.26$ $\mathbf{53.53}$ $8.44$ -
@@
Recently, open NER [10, 11, 12] has gained significant attention for its ability to generalize to new and unseen entities, offering a flexible alternative to traditional NER systems that are restricted by fixed entity types. Zhou et al. (2023) [10] introduce a targeted distillation method, where student models are trained through instruction tuning. Zaratiana et al. (2023) [11] and Bogdanov et al. (2024) [12] suggested using a tailored, small size transformer encoders, designed specifically for the NER problem, and showed that these models can outperform LLM based NER models. Sainz et al. (2023) [13] propose a method for improving information extraction results through annotation guidelines, and Ding et al. (2024) [14] explored the importance of “negative instances,” i.e, non-entity spans, for improving contextual information and label boundaries. However, integrating open type NER directly into ASR systems remains an unexplored area, as current end-to-end NER models can only operate with a fixed predefined entity set.
@@
In this paper, we propose WhisperNER, a unified model that simultaneously performs speech transcription and named entity recognition, based on the Whisper ASR model [1]. By integrating NER directly into the ASR process, WhisperNER eliminates the need for separate ASR and NER components, effectively reducing the risk of error propagation. WhisperNER closes the gap between SLU based NER approaches and the recent NLP-based open-type NER methods. By supporting open-type NER, WhisperNER can identify a broad range of entities at inference, making entity extraction from speech more accurate and flexible. To build WhisperNER, we leverage a large-scale synthetic dataset released by Bogdanov et al. (2024)  [12]. We augment a subset of the textual open NER dataset with corresponding synthetic audio samples. The final dataset contains 350350K samples and 1.81.8M unique entity types. Our approach draws motivation from recent approaches, which incorporate prior information and context into the decoding process by prompting the model on domain information [15] or keyword phrases [16, 17]. We train WhisperNER by prompting the model with NER type, for the task of joint transcription and NER tagging as depicted in Figure 1. At inference time, WhisperNER can generalize to novel entity types not observed during training.
@@
This paper makes the following contributions: (i) propose WhisperNER, a novel model for joint open NER and ASR; (ii) introducing new speech and text NER and open-NER benchmarks; (iii) conduct extensive experiments to demonstrate the superiority of WhisperNER over existing methods. To encourage future research and reproducibility, we make our source code, datasets, and models publicly available.
@@ II Our Method
In this section, we describe our approach, WhisperNER, which extends the Whisper ASR model [1] to allow for joint open NER and speech recognition. The model is designed to output a tagged transcription sequence given an audio input, effectively performing both tasks simultaneously.
Figure 2: Entity bias value: The effect of adding a bias to the logits of the entity start token (<). This simple approach allows the user significant control over the precision-recall tradeoff.
@@ II-A Model Architecture and Training
Denote by $\mathbf{x}$ the input audio features, which are fed into the encoder. The encoder produces a sequence of hidden states $\mathbf{h}=\text{Encoder}(\mathbf{x})$, which are then used to condition the decoding process. In our approach, we further condition the decoding process on a set of entity tags $\mathbf{t}=[t_{1},t_{2},\ldots,t_{k}]$, where each $t_{i}$ represents a specific entity type such as occupation, celestial-body, etc.
The decoder generates each token $y_{t}$ based on the previous tokens and the encoder’s hidden states,
@@
Overall, the output of the decoder is a sequence of tokens $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$, comprising both the transcribed text and the corresponding entity tags, as illustrated in Figure 1.
@@
To mark the entities in the output sequence, we explore two approaches. The first approach uses the BIO tagging scheme, similar to the approach used in [14], i.e., The(O) astronaut(B-occupation) explored(O).... The second marks entities with start and end markers as follow $t_{i}$$y_{i_{1}}...y_{i_{n}}$$t_{i}$>>..., for example, The <occupation>astronaut<occupation>> explored (see Figure 1). While the former may induce greater supervision and improve generalization [14], this approach increases the length of the generated sequence, hence it is potentially less computationally and time efficient.
Our prompting based training procedure, allows WhisperNER to handle both a predefined set of entities and open-type NER tasks. By exposing the model to diverse entity types during training, we ensure that it can generalize well to new and unseen entities during inference, making it highly adaptable to a wide range of applications.
@@
II-B Precision-Recall Control via Logit Modification
In standard discriminative NER models, entity probabilities are explicitly computed, allowing precision-recall trade-offs to be adjusted by varying the decision threshold [11, 18, 19, 20]. However, in generative models like LLMs [21, 22, 23] or end-to-end approaches with a generative decoder [6, 9, 24], as well as in WhisperNER, probability-based thresholding is not directly applicable due to the model’s autoregressive nature and the implicit generation of entity spans.
To address this, we propose a precision-recall control method by adjusting the logits of the entity start token, <, during inference. Specifically, we introduce a tunable bias term to the raw logits (pre-softmax) of the entity start token, shifting their relative confidence levels. A positive bias enhances entity extraction, improving recall, while a negative bias makes predictions more conservative, favoring higher precision. This straightforward yet effective approach allows users to control the precision-recall tradeoff without modifying the model architecture or requiring retraining. We evaluate this method in Section III-B and show it allows significant control over the precision-recall tradeoff curve, which may be essential for real-world applications  [25, 26].
II-C Negative Sampling and Entity Type Dropout
Empirically, we found that prompting the model with entity labels that are absent from the speech utterance (negative entity labels) significantly improves its generalization performance. This aligns with previous findings in the literature on open-type NER [11]. The negative set of entity labels was taken from a randomly selected example. Additionally, we applied random entity type dropout during training, as inspired by Sainz et al. (2023) [13]. This involves randomly eliminating a subset of entity types from the prompt and output sequence, compelling the model to focus on the remaining defined entity types. We found this approach to reduce entity hallucination risks during inference – where the model outputs entities not presented in the input prompt. Also, we incorporate random shuffling of entities in the input prompt to further promote generalization across diverse input sequences. For a detailed analysis of these methodologies and their impacts, see Section III-C.
@@ III Experiments
In this section, we evaluate WhisperNER against natural pipeline-based baselines using open-type NER speech datasets and common NER. To construct a dataset for evaluation, we either augment common NER text datasets with corresponding synthetic speech utterances or annotating the transcription of speech dataset with open-type NER labels. To support future research on joint ASR and NER, we will release all datasets.
@@
Datasets.  For the open type NER benchmarks we utilize three frequently used speech datasets with diverse linguistic and contextual properties. These datasets include VoxPopuli [27], LibriSpeech [28], and Fleurs [29], each offering unique challenges and opportunities for evaluating NER models, particularly in open-type NER contexts. For the multilingual datasets, we focus specifically on the English subset of the data. As these ASR datasets do not include NER labels, we adopt an LLM-based annotation approach, in line with the method described in [12]. See [12] for full details of the LLM-based annotation method. This annotation scheme results in the following ASR-NER datasets: (i) VoxPopuli-NER: This dataset comprises $879$ samples. It is annotated with $2469$ unique entity types, providing a comprehensive basis for political and geographical entity recognition such as: Political Role, Demographic Group, Political Institution, Political Strategy and Social Phenomenon; (ii) LibriSpeech-NER: Based on the LibriSpeech-clean dataset, with $1604$ samples derived from audio book readings, includes annotations for $3674$ unique entity types. This dataset is particularly valuable for testing the model’s ability to identify personal names and locations within narrative contexts; (iii) Fleurs-NER: Consists of $441$ samples, annotated for $1440$ unique entity types.
@@
In addition, we utilize commonly used, textual NER benchmarks [11, 12, 13, 14], namely MIT-Movie [30] and MIT-Restaurant [30]. These datasets consists of $\sim 14$K and $\sim 10$K, with $12$ and $8$ entity types, respectively. To enhance these text-based datasets for speech recognition, we generated corresponding audio data using the AWS Polly text-to-speech (TTS) service.
@@
Baselines.  We evaluate our WhisperNER againt natural pipeline-based baselines, in which we initially employ a Whisper large-v2 model for speech-to-text transcription, followed by a NER model. We constrain our evaluation to recent small-scale NER models (up to 0.5 billion parameters) to ensure competitiveness and maintain parity in terms of the number of parameters and computational requirements, aligning with our end-to-end approach. Importantly, current end-to-end models support only a fixed and predefined set of entity types, making direct comparison on the open-type NER benchmarks unsuitable [6, 9, 7, 24]. We evaluate the following approaches: (i) GLiNER [11]: a transformer encoder based architecture with zero-shot NER capabilities. We use the large variant, GLiNER-L; (ii) NuNER [12]: since the original NuNER model is not suitable for the zero-shot NER task, we utilize a GLiNER based architecture model, trained with the synthetic dataset of [12], named NuNER zero-span.; (iii) GNER [14]: Different from the previous two benchmarks, this approach uses generative models specifically adopted for the NER task. We use the flan-t5-base variant; (iv) WhisperNER-BIO: Our approach with complete BIO annotations; (v) WhisperNER: Our approach with NER span boundary annotations. (Figure 1).
@@
TABLE II: Effect of negative sampling approach: We evaluate WhisperNER models trained using the NuNER dataset for 25K steps, on the zero-shot task over the MIT NER Benchmarks.
Training details.  The WhisperNER models are trained using a subset of 350350K samples from the dataset proposed in [12]. For each example, we generate a corresponding synthetic audio sample. Unless stated otherwise, we train the WhisperNER models for 250250K steps, with learning rate of 1e61e-6 and linear decay learning rate scheduler. During training, we freeze the Whisper encoder and only modify the decoder.